OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0. quantity of negative 9 plus or minus 3 square root of 5 all over 2 quantity of 9 plus or minus 3 square root of 5 all over 2 quantity of negative 9 plus or minus square root of 117 all over 2 quantity of 9 plus or minus square root of 117 all over 2

OpenStudy (anonymous):

@Luigi0210 please help

OpenStudy (luigi0210):

Looks like you'll have to use the quadratic formula: \[\LARGE \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

can you show the steps

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Simple, plug in values and simplify: \[\LARGE x^2+9x+9\] a=1, b=9, c=9 \[\LARGE ax^2+bx+c\]

OpenStudy (anonymous):

(−9±35√)/2 Source: http://openstudy.com/study#/updates/52841feae4b0218bc6514d3d

OpenStudy (anonymous):

so it b? @Luigi0210

OpenStudy (anonymous):

no a

OpenStudy (anonymous):

\[-9±3\sqrt{5}2\]

OpenStudy (anonymous):

Yes it would be A!

OpenStudy (anonymous):

:)